A point source emits 86.9 W of sound isotropically. A small microphone intercepts the sound in an area of 0.231 cm2, 270 m from the source. Calculate (a) the sound intensity there and (b) the power intercepted by the microphone.
Answer #1

a) I = 86.9/(4*pi*270^2)=9.49E-5


b) P = I Amicrophone = 9.49E-5*0.231E-4=2.19E-9